# Steering Embedding Models with Geometric Rotation: Mapping Semantic Relationships Across Languages and Models

### v1

@@ Abstract
Understanding how language and embedding models encode semantic relationships is fundamental to model interpretability and control. While early word embeddings exhibited intuitive vector arithmetic (”king” - ”man” + ”woman” = ”queen”), modern high-dimensional text representations lack straightforward interpretable geometric properties. We introduce Rotor-Invariant Shift Estimation (RISE), a geometric approach that represents semantic transformations as consistent rotational operations in embedding space, leveraging the manifold structure of modern language representations. RISE operations have the ability to operate across both languages and models with high transfer of performance, suggesting the existence of analogous cross-lingual geometric structure. We evaluate RISE across three embedding models, three datasets, and seven morphologically diverse languages in five major language groups. Our results demonstrate that RISE consistently maps discourse-level semantic transformations with distinct grammatical features (e.g., negation and conditionality) across languages and models. This work provides the first systematic demonstration that discourse-level semantic transformations correspond to consistent geometric operations in multilingual embedding spaces, empirically supporting the Linear Representation Hypothesis at the sentence level.
@@ 1 Introduction
Understanding how contemporary language models encode and manipulate semantic knowledge has become a central challenge in deep learning interpretability. The ability to interpret (probe) and control (steer) these internal representations is fundamental to developing trustworthy, safe AI systems. In word2vec (Mikolov et al., 2013a) and similar models, semantic relationships could be captured with simple vector arithmetic in the embedding space (i.e. the famous “king” - “man” + “woman” = “queen” analogy). This geometric transparency offered both interpretability and controllability, enabling researchers to navigate semantic space through intuitive mathematical operations.
@@
However, this clarity has largely disappeared in modern transformer-based language models. While large language models (LLMs) have achieved remarkable performance across diverse language tasks (Achiam et al., 2023; Touvron et al., 2023), their internal workings remain largely opaque (Elhage et al., 2022; Rogers et al., 2021), limiting our ability to understand, predict, and control their behavior in critical applications. Unlike the interpretable directions found in static word embeddings, the geometry of modern text representations lacks the same straightforward correspondence to semantic operations. This opacity poses significant challenges for understanding how these models organize linguistic knowledge and limits our ability to control their behavior in principled ways.
@@
The central challenge lies in identifying which geometric operations correspond to meaningful semantic transformations in these complex representation spaces. Current approaches often rely on task-specific probes (Rogers et al., 2021; Hewitt & Manning, 2019; Alain & Bengio, 2017) or steering vectors (Zou et al., 2023; Wang et al., 2023; Turner et al., 2023; Merullo et al., 2023; Trager et al., 2023), but lack generalizable frameworks for systematically mapping semantic relationships to geometric structure. Without such principled methods, we cannot determine whether the geometric regularities that made static word embeddings interpretable persist in modern language or embedding models, albeit in more complex forms.
@@
We address this gap by introducing Rotor-Invariant Shift Estimation (RISE), a geometric approach that represents semantic transformations as consistent rotational operations in embedding space, leveraging the manifold structure of modern language representations. RISE is a rotor-based alignment method that identifies cross-lingual and cross-model geometric transformations. Specifically, we demonstrate how RISE identifies and generalizes three discourse-level semantic changes (negation, conditionality, and politeness) across seven morphologically distinct languages. The goal of this study is to develop a framework for identifying discourse-level semantic changes that correspond to consistent geometric transformations, and determine how well these transformations can be mapped across model architectures and typologically diverse languages. Our approach treats semantic transformations as rotations on the unit hypersphere where sentence embeddings reside, enabling us to align different linguistic contexts into a common geometric framework. This paper presents evidence that certain semantic transformations exhibit generalizable geometric structure while others vary based on context-dependence, extending the linear representation hypothesis to cross-lingual discourse. We demonstrate this through systematic experiments across three models and seven languages, revealing that negation, conditionality, and politeness transformations can be captured as consistent rotational operations.
@@ 2.1 Linear Representation Hypothesis
The linear representation hypothesis (LRH), or linear subspace hypothesis, has emerged as a promising theory for bridging the interpretability gap for embeddings (Mikolov et al., 2013b; Levy & Goldberg, 2014; Bolukbasi et al., 2016; Ethayarajh, 2019; Park et al., 2024; 2025). The LRH posits that semantic concepts are encoded as linear structures within embedding spaces, meaning linear algebraic operations can be used for interpretation and control (e.g., ”king” - ”man” + ”woman” = ”queen” presented by Mikolov et al. (2013b)). Park et al. (2024) formalized the LRH by unifying three distinct notions of linearity that had developed independently across the literature:
@@
    word2vec-like embedding differences (Arora et al., 2016; Mimno & Thompson, 2017; Ethayarajh et al., 2018; Reif et al., 2019; Li et al., 2020; Hewitt & Manning, 2019; Chen et al., 2021; Chang et al., 2022; Jiang et al., 2023),
@@
    logistic probing (Alain & Bengio, 2017; Kim et al., 2018; nostalgebraist, 2020; Belinkov, 2022; Li et al., 2022; Geva et al., 2022; Nanda et al., 2023), and
@@
    steering vectors (Wang et al., 2023; Turner et al., 2023; Merullo et al., 2023; Trager et al., 2023).
Park et al. (2024) theoretical framework addresses a critical gap by synthesizing the first formalization of what “linear representation” means. However, while the LRH has been validated primarily within individual languages, there remains a significant gap in understanding how semantic transformations generalize across linguistic contexts. Most existing work examines static concept encodings (Park et al., 2025; 2024) rather than dynamic semantic transformations that reflect real-world language use. Our work is the first to extend the LRH to multilingual contexts and embedding models, though the linear representations we consider are not Euclidean lines but geodesic flows.
@@ 2.2 Linear & Geometric Representation Techniques
The geometric foundations established by Park et al. (2024) are crucial for understanding when and why linear algebraic operations succeed in capturing semantic relationships. With traditional Euclidean geometry, it is hard to accept that arbitrary dot products or cosine similarities have semantic meaning. Moreover, Park et al. (2024) demonstrated that the choice of inner product fundamentally determines the interpretability of geometric operations, providing principled foundations for representation analysis. Our work builds directly on recent advances in understanding linear representations in language models (Park et al., 2024; Li et al., 2023). RISE implements a technique that respects semantic structure, similar to the geometric framework developed by Park et al. (2024). While previous work focused primarily on categorical concepts and word-level transformations, RISE extends our understanding to sentence-level, discourse-level transformations and cross-lingual analysis with seven morphologically diverse languages.
@@ 2.2.1 Steering Vectors & Embedding Models
The practical applications of linear representation theory have been explored through steering vector techniques. Turner et al. (2023), Liu et al. (2023), and Zou et al. (2023) demonstrated that targeted modifications to internal, latent space representations can systematically alter model behavior without parameter updates. The majority of steering vector research (Im & Li, 2025; Rimsky et al., 2023; Zou et al., 2023; Li et al., 2023) is connected to activation steering, only investigating the impact of steering vectors in the activation, hidden, and/or latent layer of an LLM. Recently, Pham & Nguyen (2024) introduced Householder Pseudo-Rotation (HPR), which addresses activation norm consistency issues in LLM behavioral modification through direction-magnitude decomposition and pseudo-rotational transformations. Building on the insight that geometric approaches outperform additive methods, our work extends geometric reasoning to semantic transformations in embedding space through Riemannian operations. To our knowledge, there is no work investigating the application of steering vectors to embedding models – only completion models. This study contributes theoretical insights to the usefulness of steering vectors for embedding model interpretation and control.
2.3 Challenges in Generalization and Reliability
Current knowledge about the generalization properties of linear representations reveals significant limitations. The taxonomy of generalization research in natural language processing (NLP) (Hupkes et al., 2023) provides a framework for evaluating robustness, but systematic applications to representation-based techniques (i.e., steering, probing, or embedding manipulation) have been limited. Recent empirical studies have revealed that steering vector effectiveness varies substantially across different inputs and contexts (Tan et al., 2024). Secondly, the relationship between local and global linearity represents a particularly critical gap in current understanding. There have been numerous demonstrations of local linear behavior within specific domains or prompt formats, but achieving global linearity (generalizable to multiple model architectures with different pre-training) as required by strong versions of the LRH, remains challenging. While many studies demonstrate impressive results in controlled settings, they often fail to address the robustness needed in practical applications. This study contributes to the literature gap by presenting a robust framework for geometrically identifying discourse-level semantic changes across typologically diverse languages and model architectures.
@@ 3 Theoretical Motivation
The limitations identified above point toward a fundamental theoretical challenge: existing approaches operate in Euclidean space while modern embeddings live on curved manifolds. This geometric mismatch may explain why steering vectors show inconsistent cross-context performance and why linear methods struggle with robust generalization. We hypothesize that discourse-level semantic transformations correspond to intrinsic geometric operations on the embedding manifold, rather than fixed directions derived from Euclidean computations. If semantic transformations can be characterized as consistent rotational operations on the unit hypersphere where embeddings reside, this would provide both theoretical grounding for the Linear Representation Hypothesis in curved spaces (through geodesics) and practical tools for cross-lingual semantic control. Testing this hypothesis requires robust evaluation across diverse languages and embedding architectures to determine whether geometric consistency reflects universal semantic properties or model-specific artifacts.
@@ 4 Rotor-Invariant Shift Estimation (RISE)
Modern sentence embeddings from multilingual encoders reside approximately on a unit hypersphere in high-dimensional space when the training objective enforces or fixes the $\ell_{2}$-norm constraints (Hirota et al., 2020), the embeddings are normalized to unit length (Reimers & Gurevych, 2019), or the model is designed to produce isotropic embeddings (Li et al., 2020; Ethayarajh, 2019). Local semantic transformations (e.g., negation, politeness, conditionality) can be understood as rotational displacements on this sphere. The key insight is that these displacements can be interpreted and controlled by aligning different contexts to a common geometric frame.
@@
where $\log_{n}(v)$ denotes the Riemannian logarithm that computes the tangent vector from $n$ to $v$ on the hypersphere, and $R(n)$ aligns the tangent vector to the canonical reference direction. Normalized embeddings reside on a unit hypersphere, where geodesics define the shortest paths between points, preserving the manifold’s intrinsic geometry rather than imposing Euclidean distance measures. These geodesic paths represent the natural notion of “line” in the embedding space, as they define the shortest distance between two points on the surface. By working with geodesics, we ensure our semantic transformations are consistent with the manifold structure. To “flatten” out the curved arc to a straight vector, the Riemannian logarithmic map $\log_{n}(v)$ produces the vector from $n$ to $v$ on a tangent plane at $n$. By operating within the tangent space at $n$, geodesic differences can be treated as ordinary vectors.
@@ 4.1 The Rotor Alignment Algorithm
RISE proceeds in three steps:
@@
$R(n^{\ast})\top\vec{p}$ rotates $\vec{p}$ into the tangent space at $n^{\ast}$ . Then the Riemannian exponential $\exp_{n^{\ast}}()$ takes the tangent vector $\vec{p}$ and moves along the geodesic starting at $n^{\ast}$. The vector direction is which geodesic to follow and the length is how far along that arc to go (in radians).
@@ 5 Experimental Design
5.1 Discourse-level Semantic Changes & Language Selection
We focus on three discourse-level semantic transformations that vary in their context-dependence:
@@ 5.1 Discourse-level Semantic Changes & Language Selection
We selected seven morphologically diverse languages to ensure broad coverage of morphological, syntactic phenomena, and resource levels: English, Spanish, Japanese, Tamil, Thai, Arabic, and Zulu. This selection spans multiple language families (Indo-European, Sino-Tibetan, Dravidian, Afroasiatic, Niger-Congo) and different morphological types (analytic, agglutinative, fusional). The languages also represent different levels of language model availability and resources. The diversity is crucial because different languages realize semantic transformations through distinct linguistic mechanisms. For instance, negation might be expressed through: (1) Particles (i.e. English “not”); (2) Affixes (i.e. Tamil verb-internal negation, Japanese “nai”); and (3) Auxiliary constructions (various languages). By testing across this range, we can determine whether geometric consistency reflects universal semantic properties or is merely an artifact of particular linguistic structures.
5.2 Datasets & Embedding Models
@@
We use three datasets and three models for evaluation. We synthetically generated one dataset, referred to as the Synthetic Multilingual dataset, and used two open-source datasets: The Benchmark of Linguistic Minimal Pairs (BLiMP) (Warstadt et al., 2020) and Sentences Involving Compositional Knowledge (SICK) (Marelli et al., 2014). For each language-transformation combination in the Synthetic Multilingual dataset, we generated 1,000 neutral-transformed sentence pairs using GPT-4.5 with carefully controlled prompts (see Appendix C). To ensure robust analysis, we implemented several diversity controls (see Appendix D).
@@
We compare three multilingual embedding models: OpenAI’s text-embedding-3-large (OpenAI, 2024), Beijing Academy of AI’s bge-m3 (Chen et al., 2024), and Google’s mBERT (Devlin et al., 2019). The text-embedding-3-large model produces 3072-dimensional vectors, bge-m3 produces 1024-dimensional vectors, and mBERT produces 768-dimensional vectors. All selected models produce constant-length embeddings that reside on a hypersphere making them suitable for our geometric analysis. This dimensional diversity allows us to test whether RISE effectiveness depends on embedding dimensionality. We calculate a rotor alignment score where the scores represent mean cosine similarity between predicted embedding vectors and the semantically transformed pair on held-out test sets, with higher values indicating more consistent geometric structure.

### v2

# Mapping Semantic & Syntactic Relationships with Geometric Rotation

@@ Abstract
Understanding how language and embedding models encode semantic relationships is fundamental to model interpretability. While early word embeddings exhibited intuitive vector arithmetic (“king” - “man” + “woman” = “queen”), modern high-dimensional text representations lack straightforward interpretable geometric properties. We introduce Rotor-Invariant Shift Estimation (RISE), a geometric approach that represents semantic-syntactic transformations as consistent rotational operations in embedding space, leveraging the manifold structure of modern language representations. RISE operations have the ability to operate across both languages and models without reducing performance, suggesting the existence of analogous cross-lingual geometric structure. We compare and evaluate RISE using two baseline methods, three embedding models, three datasets, and seven morphologically diverse languages in five major language groups. Our results demonstrate that RISE consistently maps discourse-level semantic-syntactic transformations with distinct grammatical features (e.g., negation and conditionality) across languages and models. This work provides the first demonstration that discourse-level semantic-syntactic transformations correspond to consistent geometric operations in multilingual embedding spaces, empirically supporting the linear representation hypothesis at the sentence level.
@@ 1 Introduction
Understanding how contemporary language models encode and manipulate semantic knowledge has become a central challenge in deep learning interpretability. The ability to interpret (probe) and control (steer) these internal representations is fundamental to developing trustworthy, safe AI systems. In word2vec (Mikolov et al., 2013a) and similar models, semantic relationships could be captured with simple vector arithmetic in the embedding space (i.e. the famous “king” - “man” + “woman” = “queen” analogy). This linear transparency offered both interpretability and controllability, enabling researchers to navigate semantic space through intuitive mathematical operations.
@@
However, this clarity has largely disappeared in modern transformer-based language models. While large language models (LLMs) have achieved remarkable performance across diverse language tasks (Achiam et al., 2023; Touvron et al., 2023), their internal workings remain largely opaque (Elhage et al., 2022; Rogers et al., 2021), limiting our ability to understand, predict, and control their behavior in critical applications. Unlike the interpretable, linear directions found in static word embeddings, the geometry of modern text representations lacks the same straightforward correspondence to semantic operations. This opacity poses significant challenges for understanding how these models organize linguistic knowledge and limits our ability to interpret their behavior in principled ways.
@@
The central challenge lies in identifying which geometric operations correspond to meaningful semantic transformations in these complex representation spaces. Current approaches often rely on task-specific probes (Rogers et al., 2021; Hewitt and Manning, 2019; Alain and Bengio, 2017) or steering vectors (Zou et al., 2023; Wang et al., 2023; Turner et al., 2023; Merullo et al., 2024; Trager et al., 2023), but lack generalizable frameworks for systematically mapping semantic relationships to geometric structure. Without such principled methods, we cannot determine whether the geometric regularities that made static word embeddings interpretable persist in modern language or embedding models, albeit in more complex forms.
@@
We address this gap by introducing Rotor-Invariant Shift Estimation (RISE), a geometric approach that represents semantic-syntactic transformations as consistent rotational operations in embedding space, leveraging the manifold structure of modern language representations. RISE is a rotor-based alignment method that identifies cross-lingual and cross-model geometric transformations. Specifically, we demonstrate how RISE identifies three discourse-level semantic-syntactic changes (negation, conditionality, and politeness) across seven morphologically distinct languages and generalizes across three different embedding model architectures. The goal of this study is to develop a framework for identifying discourse-level semantic-syntactic changes that correspond to consistent geometric transformations, and determine how well these transformations can be cross-lingually mapped across model architectures. Our approach treats semantic-syntactic transformations as rotations on the unit hypersphere, where sentence embeddings reside, enabling us to align different linguistic contexts into a common geometric framework. This paper presents evidence that certain semantic-syntactic transformations exhibit generalizable geometric structure while others vary based on context-dependence, extending the linear representation hypothesis to cross-lingual discourse. We demonstrate this through empirical experiments across two baselines, three models, and seven languages – revealing that negation, conditionality, and politeness transformations can be captured as consistent rotational operations  [^1] The link to our GitHub repository is https://github.com/fuelix/RISE-steering..
@@ 2.1 Linear Representation Hypothesis
The linear representation hypothesis (LRH), or linear subspace hypothesis, has emerged as a promising theory for bridging the interpretability gap for embeddings (Mikolov et al., 2013b; Levy and Goldberg, 2014; Bolukbasi et al., 2016; Ethayarajh, 2019; Park et al., 2024; 2025). The LRH posits that semantic concepts are encoded as linear structures within embedding spaces, meaning linear algebraic operations can be used for interpretation and control (e.g., “king” - “man” + “woman” = “queen” presented by Mikolov et al. (2013b)). Park et al. (2024) formalized the LRH by unifying three distinct notions of linearity that had developed independently across the literature:
@@
    word2vec-like embedding differences (Arora et al., 2016; Mimno and Thompson, 2017; Ethayarajh et al., 2019; Reif et al., 2019; Li et al., 2020; Hewitt and Manning, 2019; Chen et al., 2021; Chang et al., 2022; Jiang et al., 2023; Mitchell and Lapata, 2008; Baroni and Zamparelli, 2010)
@@
    logistic probing (Alain and Bengio, 2017; Kim et al., 2018; Belinkov, 2022; Li et al., 2022; Geva et al., 2022; Nanda et al., 2023)
@@
    steering vectors (Wang et al., 2023; Turner et al., 2023; Merullo et al., 2024; Trager et al., 2023)
Park et al.’s theoretical framework addresses a critical gap by synthesizing the first formalization of what “linear representation” means (Park et al., 2024). However, while the LRH has been validated primarily within individual languages (monolingually), there remains a significant gap in understanding how semantic-syntactic transformations generalize across linguistic contexts (cross-lingually). Most existing work examines static concept encodings (Park et al., 2025; 2024) rather than dynamic semantic-syntactic transformations that reflect real-world language use. Our work is the first to extend the LRH to multilingual contexts and embedding models. Although, the linear representations we consider are geodesic arcs and not Euclidean lines.
@@ 2.2 Linear & Geometric Representation Techniques
The geometric foundations established by Park et al. (2024) are crucial for understanding when and why linear algebraic operations succeed in capturing semantic relationships. With traditional Euclidean geometry, it is hard to accept that arbitrary dot products or cosine similarities have semantic meaning. Moreover, Park et al. (2024) demonstrated that the choice of inner product fundamentally determines the interpretability of geometric operations, providing principled foundations for representation analysis. Our work builds directly on recent advances in understanding linear representations in language models (Park et al., 2024; Li et al., 2023). RISE implements a technique that respects semantic structure, similar to the geometric framework developed by Park et al. (2024). While previous work focused primarily on categorical concepts and word-level transformations, RISE extends our understanding to sentence-level, discourse-level transformations through cross-lingual and cross-model analysis using seven morphologically diverse languages.
@@ 2.2.1 Steering Vectors & Embedding Models
The practical applications of linear representation theory have been explored through steering vector techniques. Turner et al. (2023), Liu et al. (2024), and Zou et al. (2023) demonstrated that targeted modifications to internal, latent space representations can systematically alter model behavior without parameter updates. The majority of steering vector research (Im and Li, 2025; Rimsky et al., 2023; Zou et al., 2023; Li et al., 2023) is connected to activation steering, only investigating the impact of steering vectors in the activation, hidden, and/or latent layer of an LLM. Recently, Pham and Nguyen (2024) introduced Householder Pseudo-Rotation (HPR), which addresses activation norm consistency issues in LLM behavioral modification through direction-magnitude decomposition and pseudo-rotational transformations. Building on the insight that geometric approaches outperform additive methods, our work extends geometric reasoning to semantic transformations in embedding space through Riemannian operations. To our knowledge, there is no work investigating the application of steering vectors to embedding models – only completion models. This study extends steering principles to embedding models on manifolds, not activation-level steering.
2.3 Generalization and Reliability Challenges
Current knowledge about the generalization properties of linear representations reveals significant limitations. The taxonomy of generalization research in natural language processing (NLP) (Hupkes et al., 2023) provides a framework for evaluating robustness, but systematic applications to representation-based techniques (i.e., steering, probing, or embedding manipulation) have been limited. Recent empirical studies have revealed that steering vector effectiveness varies substantially across different inputs and contexts (Tan et al., 2024). Secondly, the relationship between local and global linearity represents a particularly critical gap in current understanding. There have been numerous demonstrations of local linear behavior within specific domains or prompt formats, but achieving global linearity (generalizable to multiple model architectures with different pre-training as required by strong versions of the LRH) remains challenging. While many studies demonstrate impressive results in controlled settings, they often fail to address the robustness needed in practical applications. This study contributes to the literature gap by presenting a robust framework for geometrically identifying discourse-level semantic-syntactic changes across typologically diverse languages and model architectures.
@@ 3 Theoretical Motivation
The limitations identified in the related literature point toward a fundamental, theoretical challenge: existing approaches operate in Euclidean/linear space while modern embeddings live on curved manifolds (spherical space). This geometric mismatch may explain why steering vector research shows inconsistent cross-context performance and why linear methods struggle with robust generalization. We hypothesize that discourse-level semantic-syntactic transformations correspond to intrinsic geometric operations on the embedding manifold, rather than fixed directions derived from Euclidean computations. If semantic transformations can be characterized as consistent rotational operations on the unit hypersphere where embeddings reside, this would provide theoretical support for the extension of the Linear Representation Hypothesis in curved spaces (through geodesics) and cross-lingual interpretability. Testing this hypothesis requires robust evaluation across diverse languages and embedding architectures to determine whether geometric consistency reflects universal semantic properties or model-specific artifacts.
@@ 4 Rotor-Invariant Shift Estimation (RISE)
Modern sentence embeddings from multilingual encoders reside approximately on a unit hypersphere in high-dimensional space when the training objective enforces or fixes the $\ell_{2}$-norm constraints (Hirota et al., 2020), the embeddings are normalized to unit length (Reimers and Gurevych, 2019), or the model is designed to produce isotropic embeddings (Li et al., 2020; Ethayarajh, 2019). Local semantic transformations (e.g., negation, politeness, conditionality) can be understood as rotational displacements on this sphere. The key insight is that these displacements can be interpreted by aligning different contexts to a common geometric frame.
@@
where $\log_{n}(v)$ denotes the Riemannian logarithm that computes the tangent vector from $n$ to $v$ on the hypersphere, and $R(n)$ aligns the tangent vector to the canonical reference direction. Normalized embeddings reside on a unit hypersphere, where geodesics define the shortest paths between points, preserving the manifold’s intrinsic geometry rather than imposing Euclidean distance measures. These geodesic paths represent the natural notion of a “line” in the embedding space, as they define the shortest distance between two points on the surface. By working with geodesics, we ensure our semantic transformations are consistent with the manifold structure. To “flatten” out the curved arc to a straight vector, the Riemannian logarithmic map $\log_{n}(v)$ produces the vector from $n$ to $v$ on a tangent plane at $n$. By operating within the tangent space at $n$, geodesic differences can be treated as ordinary vectors.
@@ 4.1 The Rotor Alignment Algorithm
RISE proceeds in three steps illustrated in Figure 1:
Figure 1: RISE step-by-step illustration.
@@
$R(n^{\ast})\top\vec{p}$ rotates $\vec{p}$ into the tangent space at $n^{\ast}$ . Then the Riemannian exponential $\exp_{n^{\ast}}(\vec{p})$ takes the tangent vector $\vec{p}$ and moves along the geodesic starting at $n^{\ast}$. The vector direction is which geodesic to follow and the length is how far along that arc to go (in radians).
@@ 5 Experimental Design
5.1 Discourse-level Semantic-Syntactic Changes & Language Selection
We focus on three discourse-level semantic-syntactic transformations that vary in their context-dependence:
@@ 5.1 Discourse-level Semantic-Syntactic Changes & Language Selection
We selected seven morphologically diverse languages to ensure broad coverage of morphological, syntactic phenomena, and resource levels: English, Spanish, Japanese, Tamil, Thai, Arabic, and Zulu. This selection spans multiple language families (Indo-European, Sino-Tibetan, Dravidian, Afroasiatic, Niger-Congo) and different morphological types (analytic, agglutinative, fusional). The languages also represent different levels of language model availability and resources. The diversity is crucial because different languages realize semantic transformations through distinct linguistic mechanisms. For instance, negation might be expressed through: (1) Particles (i.e. English “not”); (2) Affixes (i.e. Tamil verb-internal negation, Japanese “nai”); and (3) Auxiliary constructions (i.e. English “does/has not”). By testing across this range, we can determine whether geometric consistency reflects universal semantic properties or is merely an artifact of particular linguistic structures.
5.2 Datasets, Embedding Models, & Linear Baselines
@@
We use three datasets and three models for evaluation. We used two open-source, external datasets: The Benchmark of Linguistic Minimal Pairs (BLiMP) (Warstadt et al., 2020) and Sentences Involving Compositional Knowledge (SICK) (Marelli et al., 2014), and synthetically generated one dataset, referred to as the Synthetic Multilingual dataset. For each language-transformation combination in the Synthetic Multilingual dataset, we generated 1,000 neutral-transformed sentence pairs using GPT-4.5 with carefully controlled prompts (see Appendix D). To ensure robust analysis, we implemented several diversity controls (see Appendix E).
@@
We compare three multilingual embedding models: text-embedding-3-large (OpenAI, 2024), bge-m3  [^3] Bge-m3 should be m3 as titled in the final version of (Chen et al., 2024), but we referenced the model as bge-m3 in this paper and figures. (Chen et al., 2024), and mBERT (Devlin et al., 2019). The text-embedding-3-large model produces 3072-dimensional vectors, bge-m3 produces 1024-dimensional vectors, and mBERT produces 768-dimensional vectors. All selected models produce constant-length embeddings that reside on a hypersphere making them suitable for our geometric analysis. This dimensional diversity allows us to test whether RISE effectiveness depends on embedding dimensionality. We calculate a rotor alignment score where the scores represent mean cosine similarity between predicted embedding vectors and the semantically transformed pair on held-out test sets, with higher values indicating more consistent geometric structure. Table 1 describes how the cosine similarity scores are interpreted.
We include Mean Difference Vectors (MDV), and Procrustes alignment as baseline comparisons because they represent standard linear approaches used to model transformations in embedding spaces. MDV test whether simple difference vectors can capture semantic or cross-lingual structure, while Procrustes evaluates whether a single global rotation can align transformed embeddings. MDV is the geometrically correct analogue of the Euclidean additive method for modern spherical embeddings, providing a stronger and fairer baseline for RISE.
Cosine Similarity Range Interpretation Supporting Literature
$\geq 0.80$ Strong, consistent geometric structure Reimers and Gurevych (2019)
0.650.650.800.80 Moderate, reliable structure Mikolov et al. (2013b); Ethayarajh (2019)
0.500.500.650.65 Weak or variable structure Ethayarajh (2019); Conneau et al. (2018)
<0.30<0.30 Inconsistent or failing transformation Artetxe et al. (2018); Conneau et al. (2018)
Table 1: Interpretation of cosine similarity magnitudes used throughout this work. Higher values indicate stronger geometric consistency between predicted and target embeddings. These thresholds are stricter than prior work but remain consistent with the established interpretations in the literature.